OpenStudy (anonymous):

help quick please): A survey is taken and finds that 4,250 people in town would vote for candidate A. The survey company says their results are accurate within 2%. Write an absolute value inequality with x as the possible survey results. The quotient of a number and five is eight less than the number. What is the number?

OpenStudy (anonymous):

@agent0smith do you think you could help me with these two please?

OpenStudy (ranga):

With 2% error, x will be between: 4250 * .98 <= x <= 4250 * 1.02

OpenStudy (anonymous):

so would it be x+4250<0.002 ? @ranga

OpenStudy (anonymous):

*0.02

OpenStudy (ranga):

x is the survey result which is the number of votes candidate A gets. That number is 4250. But there is a margin of error of 2% in the survey. That means, it could be 2% more than 4250 or 2% less than 4250 2% more means multiply by 1.02. 4250 * 1.02 = 4335 2% less means multiplying by 0.98. 4250 * 0.98 = 4165 So 4165 <= x <= 4335 is the inequality.

OpenStudy (anonymous):

thanks! by any chance could you help me with the second problem? @ranga

OpenStudy (ranga):

Let the number be N. "The quotient of a number and five" means N / 5 "eight less than the number" is N - 8 N/5 = N - 8 Solve for N.

OpenStudy (ranga):

N/5 = N - 8 Multiply both sides by 5: N = 5N - 40 subtract N 0 = 4N - 40 add 40 40 = 4N divide by 4: 10 = N The number is 10.